# Two Effects, One Trigger: On the Modality Gap, Object Bias, and Information Imbalance in Contrastive Vision-Language Models

### v2

@@ Abstract
Contrastive vision-language models (VLMs), like CLIP, have gained popularity for their versatile applicability to various downstream tasks. Despite their successes in some tasks, like zero-shot object recognition, they perform surprisingly poor on other tasks, like attribute recognition. Previous work has attributed these challenges to the modality gap, a separation of image and text in the shared representation space, and to a bias towards objects over other factors, such as attributes. In this analysis paper, we investigate both phenomena thoroughly. We evaluated off-the-shelf VLMs and find that while the gap’s influence on performance is typically overshadowed by other factors, we find indications that closing the gap indeed leads to improvements. Moreover, we find that, contrary to intuition, only few embedding dimensions drive the gap and that the embedding spaces are differently organized. To allow for a clean study of object bias, we introduce a definition and a corresponding measure of it. Equipped with this tool, we find that object bias does not lead to worse performance on other concepts, such as attributes per se. However, why do both phenomena, modality gap and object bias, emerge in the first place? To answer this fundamental question and uncover some of the inner workings of contrastive VLMs, we conducted experiments that allowed us to control the amount of shared information between the modalities. These experiments revealed that the driving factor behind both the modality gap and the object bias, is an information imbalance between images and captions, and unveiled an intriguing connection between the modality gap and entropy of the logits.
@@ 1 Introduction
Finally, we investigate the fundamental question of what actually leads to the emergence of the modality gap and object bias. We identify a common cause: information imbalance. Information imbalance refers to the availability of more information in one modality than the other. For example, captions are often sparse (lossy), focusing on the most salient object(s), while images hold far more information not captured by their captions; see Figure 1. Since the caption is unknown to the image encoder, it cannot know what information of the image it needs to encode to align the image encoding with the text encoding. The best that the image encoder can do is to focus on the most salient parts of the image, i.e., the parts that are typically present in the captions. As a result, the image encoder develops a bias towards these parts. For natural language captions, these are typically object names. Besides that, the modality gap also emerges as a by-product of contrastive training under an information imbalanced regime. Here, the model trades off alignment—which is limited due to the information imbalance—with uniformity by making all images and all texts more dissimilar. We thoroughly validate our hypothesis by manipulating information imbalance in experiments on synthetic as well as real data. Our insights equip practitioners and researchers with the understanding needed to craft, for example, data filtering or caption enrichment to reduce the modality gap or the object bias.
@@ 4.1 Does a smaller modality gap lead to better performance?
Figures 3 and 1 show that a larger modality gap counter-intuitively correlates with better downstream performance. However, does this imply that a larger modality gap leads to better performance? No, as Table 1 reveals, other factors such as model or embedding size have a stronger effect on performance. It appears that the negative impact of the modality gap is overshadowed by these other factors. Indeed, when we control for these factors, we observe the expected negative correlation, i.e., a smaller gap seems to correlate with better performance (see Section D.2). Thus, the modality gap is a problem worth fighting. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Takeaway 1: A larger modality gap has mild positive correlation with downstream performance. However, there is no indication that a larger modality gap leads to a better performance; rather, it suggests the presence of common confounders (e.g., model size).
@@ 4.2 Few embedding dimensions drive the modality gap
Figure 4: Few embedding dimensions separate the modalities. Results on MS-COCO. (4(a)) We plot the absolute difference in the means of each embedding dimension between the modalities. Most dimensions have similar means for both modalities, but for some the differences are huge. (4(b)) Pairs of these high difference dimensions can perfectly separate the modalities. (4(c)) Successive removal of embedding dimensions based on the sorting of embedding dimensions from (4(a)) leads to a sharp drop, followed by a partial recovery of downstream performance, while the modality gap gradually closes (similar results for L2M). See Section D.3 for results on ImageNet.
@@ 5 Object bias is a caption presence bias
Figure 6: Object bias is caused by a per-sample caption presence bias. (6(a)) Object bias is not caused by the word frequency as attributes appear more frequently than objects. (6(b)) We trained CLIP models on MAD and changed which factor is always in the caption (i.e., caption presence bias). The models are biased towards whichever factor (e.g., color, thickness, etc.) that we made prelevant in the captions.
where $O$ denotes the object classes, $\text{sim}_{\text{obj}}^{+}$ denotes the mean similarity between the L2-normalized image embeddings that entail the same object obj, $\text{sim}_{\text{obj}}^{-}$ the mean similarity between different objects (i.e., the image does not entail the same object) and $\text{sim}_{\text{att}}^{+}$ $\text{sim}_{\text{att}}^{-}$ are similarly defined for attributes $A$. Positive values of MOAD indicate a bias towards objects, negative values a bias towards attributes, and zero no bias. Refer to Section E.1 for the detailed definition. Intuitively, MOAD measures whether the visible/present objects or attributes have a larger influence on the similarity of two images/captions. Note that MOAD is not specific to objects and attributes, but can be used also for other pairs of factors.
@@ 6.1 How does information imbalance cause the modality gap and object bias?
Figure 7 shows that both the modality gap and object bias decrease as we reduce information imbalance. This validates our hypothesis. Notably, even when the modality gap is large at model initialization, contrastive VLMs can close it in the full information setting (Figure 7(a)(I) and qualitatively in Figure 7(b)). This is aligned with our hypothesis. In contrast, the cone effect hypothesis is not consistent with our observation. Further, we find that the image encoders are more biased towards objects than the text encoders (Figure 5(a), 6(b) and 12(III-IV)). This is expected, as the text encoders know what to encode, whereas the image encoders must model the uncertainty over the captions’ content. Finally, we find that performance improves as information imbalance decreases (Figure 7(a)(III)), and that our findings are consistent across different embedding dimensions.
@@ 6.2 Is the modality gap a bug or a feature?
Following our explanation in Section 6.1 and the hypothesis that CLIP “increases the gap to adjust the entropy”, we expect low entropy and modality gap at the start of fine-tuning followed by a sharp increase. Importantly, we would expect that entropy is similar for both, “frozen temp” and “learnable temp.” As a consequence, the modality gap should be larger when the temperature is kept frozen. Indeed, we observe that ”frozen temp.” increases the modality gap significantly more to match the same entropy (Figure 8). Thus, the modality gap can be interpreted as a feature, not a bug, as it adds flexibility to the model to change the entropy.
@@ Acknowledgments
We would also like to thank Philipp Schröppel and Sudhanshu Mittal for helpful comments on the draft and Maria A. Bravo and Simon Ging for insightful discussions.
@@ Appendix A Further details for the analysis experiments of pre-trained vision-language models
Dataset details. We ran our evaluations on MS COCO (Lin et al., 2014; Chen et al., 2015), ImageNet (Russakovsky et al., 2015), MIT-States (Isola et al., 2015), and UT-Zappos (Yu & Grauman, 2014). The datasets comprise 25000 (5000 images with 5 captions each), 50000, 12995, or 2914 test samples, respectively. MS COCO and ImageNet are standard datasets for evaluation of retrieval or object recognition performance, respectively. We used the standard evaluation protocols to image-to-text retrieval performance (text-to-image retrieval yielded similar results) or compute accuracy. MIT-States consists of 245 objects and 115 adjectives (attributes), while UT-Zappos consists of 12 shoe types with 16 fine-grained states ($\sim$ attributes). For both datasets, we assume that we do not know the object of a respective image and only want to find the adjective or fine-grained state. We considered this a classification problem, following previous work (Trager et al., 2023). Note that these datasets implicitly assume that the adjectives are mutually exclusive per image. However, this may not be necessarily true, as multiple adjectives or fine-grained states may be present in the image.
@@ Appendix C Background on the CLIP loss
  • Alignment (nominator): matching pairs should be close, i.e., aligned.
@@
For multi-modal contrastive representation learning, it is popular to use a symmetric version of above InfoNCE objective (Radford et al., 2021):
@@ D.2 Further analysis on the relationship of the modality gap and downstream performance
Table 3: When controlling for the dataset choice, the modality gap correlates negative with performance (i.e., the smaller the gap the better the performance) in most cases. We report the Kendall’s $\tau$ rank correlations between RMG (top) or L2M (bottom) and the respective downstream performance. Additionally, we report the p-values of the correlations. Note, that the p-values are sometimes large due to the small number of models we are left with for each dataset (see parenthesis) and therefore results should be taken with a grain of salt.
@@ D.5 Closing the gap for ideal words
We showed that we can close the modality gap post-hoc by successive removal of embedding dimensions in Section 4.2 but did not improve performance. Similarly, Liang et al. (2022) shifted the embeddings by the modality gap vector
@@
Ideal words (Trager et al., 2023). Ideal words can be computed as follows:
@@ Results.
Appendix E Extend details for Section 5
@@ E.1 Further details on our object attribute bias metric MOAD
E.2 Additional experiment on where the object bias stems from
@@ F.1 CC12M training and information imbalance causes the modality gap experiments
Experimental setup. We largely followed the training setup of Ilharco et al. (2021). As common for CC12M, we used a ResNet-50 image encoder. We trained for 30 epochs, with each epoch having 9263104 image-text samples. We used a global batch size of 1760 and a learning rate of $31e-$ 3 with 1000 steps of warm-up. For all other settings, we used the default values of Ilharco et al. (2021).
@@
where $\mathbf{x}_{i}$ is the $i$-th L2-normalized image embedding and $\mathbf{y}_{i}$ the $i$-th text embedding. Note that in contrast to RMG, L2I misses the drop of the relative gap (approx. epoch 25), as the distance between the positives keeps increasing.
@@ F.3 Can the modality gap even appear in the global optimum?
However, what happens when we alignment is bounded, e.g., due to mismatches? For image-text pairs, this can happen if a human annotator miscaptions an image. It could also stem from differing focuses among human annotators on distinct aspects of the image. We considered the following pairs:
@@ F.4 Is the gap primarily caused by the contrastive loss instead of the information imbalance?
Training details. We followed the experimental setting of Fahim et al. (2024). We used 2048 randomly selected images from the MS COCO validation set (they used the training set). We also trained the CLIP models with a batch size of 64 for 1200 epochs. Both image encoders use the ViT-B/32 architecture with a 512-dimensional embedding dimensionality. We used a learning rate of $41e-$ 4 with AdamW (with weight decay of $0.1$, $\beta_{1}=0.9$, $\beta_{2}=0.98$, and ${\epsilon}=1e-6$) and cosine annealing learning rate scheduling. The maximum temperature was set to 100.0. We repeated trainings for five random seeds.

### v3

@@ Abstract
Contrastive vision-language models (VLMs), like CLIP, have gained popularity for their versatile applicability to various downstream tasks. Despite their successes in some tasks, like zero-shot object recognition, they perform surprisingly poor on other tasks, like attribute recognition. Previous work has attributed these challenges to the modality gap, a separation of image and text in the shared representation space, and to a bias towards objects over other factors, such as attributes. In this analysis paper, we investigate both phenomena thoroughly. We evaluated off-the-shelf VLMs and while the gap’s influence on performance is typically overshadowed by other factors, we find indications that closing the gap indeed leads to improvements. Moreover, we find that, contrary to intuition, only few embedding dimensions drive the gap and that the embedding spaces are differently organized. To allow for a clean study of object bias, we introduce a definition and a corresponding measure of it. Equipped with this tool, we find that object bias does not lead to worse performance on other concepts, such as attributes per se. However, why do both phenomena, modality gap and object bias, emerge in the first place? To answer this fundamental question and uncover some of the inner workings of contrastive VLMs, we conducted experiments that allowed us to control the amount of shared information between the modalities. These experiments revealed that the driving factor behind both the modality gap and the object bias, is an information imbalance between images and captions, and unveiled an intriguing connection between the modality gap and entropy of the logits.
@@ 1 Introduction
Finally, we investigate the fundamental question of what actually leads to the emergence of the modality gap and object bias. We identify a common cause: information imbalance. Information imbalance refers to the availability of more information in one modality than the other. For example, captions are often sparse (lossy), focusing on the most salient object(s), while images hold far more information not captured by their captions; see Figure 1. Since the caption is unknown to the image encoder, it cannot know what information of the image it needs to encode to align the image encoding with the text encoding. The best that the image encoder can do is to focus on the most salient parts of the image, i.e., the parts that are typically present in the captions. As a result, the image encoder develops a bias towards these parts. For natural language captions, these are typically object names. Besides that, the modality gap also emerges as a by-product of contrastive training under an information imbalanced regime. Here, the model trades off alignment—which is limited due to the information imbalance—with uniformity by making all images and all texts more dissimilar. We thoroughly validate our hypothesis by manipulating information imbalance in experiments on synthetic as well as real data. Our insights equip researchers with the understanding needed to craft, for example, data filtering or caption enrichment strategies to reduce the modality gap or object bias.
@@ 4.1 Does a smaller modality gap lead to better performance?
Figures 3 and 1 show that a larger modality gap counter-intuitively correlates with better downstream performance. However, does this imply that a larger modality gap leads to better performance? No, as Table 1 reveals, other factors such as model or embedding size have a stronger effect on performance. It appears that the negative impact of the modality gap is overshadowed by these other factors. Indeed, when we control for these factors, we find weak indications for the expected negative correlation for some of the pre-training datasets, i.e., a smaller gap seems to correlate with better performance (see Section D.2). Thus, the modality gap is a problem worth fighting. {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = seabornblue!10, roundcorner = 4pt, ] Takeaway 1: A larger modality gap has mild positive correlation with downstream performance. However, there is no indication that a larger modality gap leads to a better performance; rather, it suggests the presence of common confounders (e.g., model size).
@@ 4.2 Few embedding dimensions drive the modality gap
Figure 4: Few embedding dimensions separate the modalities. Results on MS-COCO. (4(a)) We plot the absolute difference in the means of each embedding dimension between the modalities. Most dimensions have similar means for both modalities, but for some the differences are huge. (4(b)) Pairs of these high difference dimensions can perfectly separate the modalities (we show the ones with largest mean for each modality). (4(c)) Successive removal of embedding dimensions based on the sorting of embedding dimensions from (4(a)) leads to a sharp drop, followed by a partial recovery of downstream performance, while the modality gap gradually closes (similar results for L2M). See Section D.3 for results on ImageNet and the plots in (4(b)) with the largest two dimensions of (4(a)).
@@ 5 Object bias is a caption presence bias
Figure 6: Object bias is caused by a per-sample caption presence bias. (6(a)) Object bias is not caused by the word frequency as attributes appear more frequently than objects. (6(b)) We trained CLIP models on MAD and changed which factor is always in the caption (i.e., caption presence bias). The models are biased towards whichever factor (e.g., color, thickness, etc.) that we made prevalent in the captions.
where $O$ denotes the object classes, $\text{sim}_{\text{obj}}^{+}$ denotes the mean similarity between the L2-normalized image embeddings that entail the same object obj, $\text{sim}_{\text{obj}}^{-}$ the mean similarity between different objects (i.e., the image does not entail the same object) and $\text{sim}_{\text{att}}^{+}$ $\text{sim}_{\text{att}}^{-}$ are similarly defined for attributes $A$. Positive values of MOAD indicate a bias towards objects, negative values a bias towards attributes, and zero no bias. Refer to Section E.1 for the detailed definition. Intuitively, MOAD measures whether the visible/present objects or attributes have a larger influence on the similarity of two images/captions. Note that MOAD is not specific to objects and attributes, we provide a generic formulation, called Bias measure via Relative Angle Change of Embeddings (BRACE) , in Section E.2.
@@ 6.1 How does information imbalance cause the modality gap and object bias?
Figure 7(a) shows that both the modality gap and object bias decrease as we reduce information imbalance. This validates our hypothesis. Notably, even when the modality gap is large at model initialization, contrastive VLMs can close it in the full information setting (Figure 7(a)(I) and qualitatively in Figure 7(b)). This is aligned with our hypothesis. In contrast, the cone effect hypothesis is not consistent with our observation. Further, we find that the image encoders are more biased towards objects than the text encoders (Figure 5(a), 6(b) and 13(III-IV)). This is expected, as the text encoders know what to encode, whereas the image encoders must model the uncertainty over the captions’ content. Finally, we find that performance improves as information imbalance decreases (Figure 7(a)(III)), and that our findings are consistent across different embedding dimensions.
@@ 6.2 Is the modality gap a bug or a feature?
Following our explanation in Section 6.1 and the hypothesis that CLIP “increases the gap to adjust the entropy”, we expect low entropy and modality gap at the start of fine-tuning followed by a sharp increase. Importantly, we would expect that entropy is similar for both, “frozen temp” and “learnable temp.” As a consequence, the modality gap should be larger when the temperature is kept frozen. Indeed, we observe that “frozen temp.” increases the modality gap significantly more to match the same entropy (Figure 8). Thus, the modality gap can be interpreted as a feature, not a bug, as it adds flexibility to the model to change the entropy.
@@ Acknowledgments
We would like to thank Philipp Schröppel and Sudhanshu Mittal for helpful comments on the draft, Maria A. Bravo and Simon Ging for insightful discussions, and Simon Ging and Elias Kempf for preparing the CC12M and CC3M datasets.
@@ Appendix A Further details for the analysis experiments of pre-trained vision-language models
Dataset details. We ran our evaluations on MS COCO (Lin et al., 2014; Chen et al., 2015), ImageNet (Russakovsky et al., 2015), MIT-States (Isola et al., 2015), and UT-Zappos (Yu & Grauman, 2014). The datasets comprise 25000 (5000 images with 5 captions each), 50000, 12995, or 2914 test samples, respectively. MS COCO and ImageNet are standard datasets for evaluation of retrieval or object recognition performance, respectively. We used the standard evaluation protocols to image-to-text retrieval performance (text-to-image retrieval yielded similar results) or computed top-1 accuracy. MIT-States consists of 245 objects and 115 adjectives (attributes), while UT-Zappos consists of 12 shoe types with 16 fine-grained states ($\sim$ attributes). For both datasets, we assume that we do not know the object of a respective image and only want to find the adjective or fine-grained state. We considered this a classification problem, following previous work (Trager et al., 2023). Note that these datasets implicitly assume that the adjectives are mutually exclusive per image. However, this may not be necessarily true, as multiple adjectives or fine-grained states may be present in the image.
@@ Appendix C Background on the CLIP loss
  • Alignment (numerator): matching pairs should be close, i.e., aligned.
@@
For multi-modal contrastive representation learning, it is popular to use a symmetric version of above objective (Radford et al., 2021):
@@ D.2 Further analysis on the relationship of the modality gap and downstream performance
Table 3: When controlling for the dataset choice, the modality gap seems to correlate negatively with performance (i.e., the smaller the gap the better the performance) in most cases. We report the Kendall’s $\tau$ rank correlations between RMG (top) or L2M (bottom) and the respective downstream performance. Additionally, we report the p-values of the correlations. Note, that the p-values are sometimes large due to the small number of models we are left with for each dataset (see parenthesis) and therefore results should be taken with a grain of salt.
@@ D.3 ImageNet results for Figure 4
Figure 12: Pairs based on the largest mean differences. In the main paper, we show the dimensions with largest means. Note that using the largest mean difference also leads to separable dimensions. Top: OpenAI’s CLIP ViT-B/16, bottom: SigLIP ViT-B/16. Left: MS COCO, right: ImageNet. In Figures 4(b) and 11(c), we selected the embedding dimension of each modality, as each of the embedding dimensions most dominate the direction of the embedding vector of each modality.
Figure 12 shows that the modalities are also well separable when we select embedding dimensions based on the largest mean differences; similar to Figures 4(b) and 11(c).
@@ D.5 Closing the gap for ideal words
We showed that we can close the modality gap post-hoc by successive removal of embedding dimensions in Section 4.2 but performance did not improve. Similarly, Liang et al. (2022) shifted the embeddings by the modality gap vector
@@
Ideal words. Ideal words (Trager et al., 2023) can be computed as follows:
@@ Results.
Appendix E Extended details for Section 5
@@ E.1 Further details on our object attribute bias metric MOAD
E.2 General formulation for Bias measure via Relative Angle Change of Embeddings (BRACE)
MOAD is an instantiation of our novel metric Bias measure via Relative Angle Change of Embeddings (BRACE). BRACE is defined as follows:
$\begin{split}\text{BRACE}_{\mathbf{\text{img}}}\coloneqq~{}&\frac{1}{2|B|}\sum\limits_{b\in B}\left(\frac{1}{N_{1}}\sum\limits_{\begin{subarray}{c}\mathbf{x}_{i},\mathbf{x}_{j}\in X_{b}\\ i\neq j\end{subarray}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}-\frac{1}{N_{2}}\sum\limits_{\mathbf{x}_{i}\in X_{b},\mathbf{x}_{j}\in X_{\neg b}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}\right)\\ -&\frac{1}{2|C|}\sum\limits_{c\in C}\left(\frac{1}{N_{3}}\sum\limits_{\begin{subarray}{c}\mathbf{x}_{i},\mathbf{x}_{j}\in X_{c}\\ i\neq j\end{subarray}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}-\frac{1}{N_{4}}\sum\limits_{\mathbf{x}_{i}\in X_{c},\mathbf{x}_{j}\in X_{\neg c}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}\right)\quad,\end{split}$ (9)
where $x_{i}$ denotes the normalized image embeddings, $N_{1},...,N_{4}$ are normalization factors given by $N_{1}=|X_{b}|^{2}-|X_{b}|$, $N_{2}=|X_{b}\times X_{\neg b}|$, $N_{3}=|X_{c}|^{2}-|X_{c}|$ and $N_{4}=|X_{c}\times X_{\neg c}|$, $X_{b},X_{\neg b},X_{c},X_{\neg c}$ are all image embeddings $\mathbf{x}$ that (not) entail some group of concepts $B$ (e.g., objects $O$) or some other group of concepts $C$ (e.g., attributes $A$), respectively.
E.3 Additional experiment on where the object bias stems from
@@ F.1 CC12M training and information imbalance causes the modality gap experiments
Experimental setup. We largely followed the training setup of Ilharco et al. (2021). As common for CC12M, we used a ResNet-50 image encoder. We trained for 30 epochs, with each epoch having 9263104 image-text samples. We used a global batch size of 1760 and a learning rate of 1e-3 with 1000 steps of warm-up. For all other settings, we used the default values of Ilharco et al. (2021).
@@
where $\mathbf{x}_{i}$ is the $i$-th L2-normalized image embedding and $\mathbf{y}_{i}$ the $i$-th text embedding. Note that in contrast to RMG, L2I misses the drop of the relative gap (approx. epoch 25 in Figure 14), as the distance between the positives keeps increasing.
(a) ViT-S vision encoder.
(b) Dropping sentences instead of words.
(c) Other dataset: CC3M.
Figure 15: Further evidence for the effect of information imbalance on the modality gap on real data. This figure supports the findings of Figures 7(c) and 14 with experiments using (15(a)) a different vision encoder backbone, (15(b)) sentence-wise information-dropping scheme, and (15(c)) training on a different dataset (note that CC3M is not a subset of CC12M). In all settings, “full” captions (i.e., less information imbalance) also leads to a smaller modality gap (RMG, L2M).
Figure 15 shows results for a different vision encoder (ViT-S in (15(a))), an alternative information-dropping scheme (sentence-dropping in (15(b))), and another dataset (CC3M in (15(c))). For all these variations, we find that the modality gap is smaller with less information imbalance (“full” caption) similar to our findings in Figures 7(c) and 14.
@@ F.3 Can the modality gap even appear in the global optimum?
However, what happens when the alignment is bounded, e.g., due to mismatches? For image-text pairs, this can happen if a human annotator miscaptions an image. It could also stem from differing focuses among human annotators on distinct aspects of the image. We considered the following pairs:
@@ F.4 Is the gap primarily caused by the contrastive loss instead of the information imbalance?
[…]
@@ F.5 Can the contrastive loss close the modality gap?
F.6 Can the modality gap be reduced in pre-trained models?
To show that the modality gap can also be reduced for pre-trained contrastive VLMs, we fine-tuned OpenAI’s CLIP ViT-B/16 (Radford et al., 2021) and SigLIP ViT-B/16 (Zhai et al., 2023) on the Densely Captioned Images (DCI)dataset (Urbanek et al., 2024). DCI consists of 7805 dense and highly-aligned image-text pairs. Thus, DCI has reduced information imbalance and, consequently, the modality gap should reduce as we fine-tune.
[…]
Results.
MS COCO
CLIP ViT-B/16 SigLIP ViT-B/16
before fine-tuning after fine-tuning before fine-tuning after fine-tuning
L2M 0.816 0.700 1.046 0.958
RMG 0.572 0.499 0.623 0.562
R@1 52.84 59.24 67.68 67.58
ImageNet
L2M 0.864 0.747 1.116 1.04
RMG 0.606 0.531 0.683 0.63
Top-1 accuracy 66.75 65.13 75.61 73.87
Table 10: Fine-tuning on the image-text pairs of DCI (Urbanek et al., 2024) reduces the modality gap (lower L2M & RMG).
Table 10 shows that training on the image-text pairs of DCI indeed reduces the modality gap, as expected by our arguments in Section 6.1. Note that fine-tuning led to slightly worse zero-shot transfer accuracy on ImageNet. This is not in conflict with our claims, as other factors such as the distributional shift between the fine-tuning and test datasets play a larger role for performance. For example, we do not expect better performance when fine-tuning on an image-text flower dataset and then evaluate on an image-text animal dataset. We suspect that DCI similarly may not be best suited for such object-centric benchmarks (ImageNet & MS COCO).
Appendix G Limitations
We validated our main hypotheses on small- (MAD, CC3M) and medium-sized (CC12M) datasets. While additional experiments on large-scale datasets like LAION would be valuable, they are beyond our computational resources, unfortunately. Nevertheless, our analysis of pre-trained large-scale CLIP models aligns with our findings on smaller datasets, suggesting consistency across scales.
Additionally, our method of reducing the information imbalance on real data (keeping only a contiguous sequence of words) may not be optimal. For instance, dropping 50% of words should, on average, remove 50% of information, but the variance may be large. A more refined method could produce even clearer results.
Lastly, the results in Tables 1 and 3 are correlation-based. Due to the small number of comparable models, the findings in Table 3 lack statistical significance. Expanding our current analysis would require prohibitively large compute resources, which we unfortunately have not.
Appendix H Connection to partial multi-label learning
Information imbalance highlights the learning difficulties that arise from incomplete or partial information. In our case, a modality gap and object bias. This issue is also present in more classical multi-label learning. Specifically, the field of partial multi-label learning (Xie & Huang, 2018; Struski et al., 2023; Hang & Zhang, 2023) handles scenarios where an unknown number of ground truth labels are missing while irrelevant or adulterated ground truth labels are provided.
In this work, we showed that CLIP-like models face analogous challenges that stem from information imbalance. For example, information is missing from the captions, or captions entail mismatches. We believe that future work could study the similarities and potentially transfer ideas from partial multi-label learning to CLIP-like models, or vice versa.